Solve the given initial-value problem. - 5y = 0, y(1) = 0, y(1) = 9 OL
Answer #1

d24 - 4 - 5y =o__y030 , (139 Auxiliary equation above differential equation Lis m²-um-550 m²-5m+m-5=0 mm-5)+1m-5)=0 (mtilmay (1 = 0 - y(1) = 9 0= ce toes g = c e + 5 Ge - 0 and ② Add equation g = Ges - - 02 Guexe - 9 fyct) = -3 et + 3 stat